2. The growth in demand for computer hardware in a college is given by the equation t Q = )025.1(200 , where t is in months.



i) Calculate the demand when t = 0, 3, 6, 10 months.



ii) Calculate the time taken for the demand to increase to 300.



iii) Graph the demand for t = 0 to 30, hence confirm the answers to (i) and (ii).

1
Expert's answer
2022-04-28T16:23:52-0400
"Q(t)=200(1.025)^t"

i)


"Q(0)=200(1.025)^0=200"

"Q(3)=200(1.025)^3=215.378"

"Q(6)=200(1.025)^6=231.939"

"Q(10)=200(1.025)^{10}=256.017"

ii)

"200(1.025)^t=300"

"t=\dfrac{\ln1.5}{\ln1.025}"

"t=16.42\ months"

iii)